OpenStudy (grayj200):

Statistics Question: The amount of money spent weekly on cleaning, maintenance, and repairs at a large restaurant was observed over a long period of time to be approximately normally distributed, with mean μ = $605 and standard deviation σ = $47? (b) How much should be budgeted for weekly repairs, cleaning, and maintenance so that the probability that the budgeted amount will be exceeded in a given week is only 0.12? (Round your answer to the nearest dollar.)

OpenStudy (grayj200):

If I knew how to put it in the calculator correctly I would have already done that, haha.

Directrix (directrix):

What is part A of this question? I am wondering if the answer to it is related to the part B question?

Directrix (directrix):

P{ (z > (x - 605)/47 } > .12 (x - 605) = approx 1.17 x - 605 = 47 * 1.17 x = 54.99 approx @grayj200 See what you think about this. I used the table of areas and not a calculator.